OpenStudy (anikate):

9y^2-(2x-y)^2 help me with this question plz. Plz explain to me I need to factorize.

OpenStudy (anikate):

hi

OpenStudy (anonymous):

\[(2x-y)^2 = (2x-y)(2x-y) =(2x)(2x) + (2x)(-y)+(-y)(2x)+(y)(y) = 4x^2 - 4xy + y^2\] \[= {9y^2 - (4x^2 -4xy + y^2)} = 9y^2 - 4x^2 + 4xy - y^2 = 8y^2 + 4xy -4x^2 = 4(2y^2+xy-x^2)\]